BlueWolf08:

Equation 1: (x – 12) = 12 Equation 2: y – 12 = 12 Solve each equation. Halp math hw

BlueWolf08:

Halp meh

mxddi3:

so, if I am correct, we need to find what x= and what y= ?

BlueWolf08:

Yeah

mxddi3:

ok uhm do your parentheses mean something, or can we just remove them? but for the y-12=12 you need to get the y on a side by itself, which means you need to add 12 to both sides. if the parentheses do not mean anything, then do the same with x.

BlueWolf08:

Usually it means do it first so I think we do that before anything else then take the parenthesis away

mxddi3:

well we do not know what the X is in this case, so we are unable to solve (x-12) only way to do so would be to solve for x first. so as i said, I think you would have (x-12)=12 then you need to solve this. and do the same for y. are you able to do this?

BlueWolf08:

okk 63

BlueWolf08:

oop not 63 just a typo

mxddi3:

that's fine. one you solve (if you need help doing so i can help you) please let me know the answer you got. you can also show your work here if you'd like

BlueWolf08:

ok

BlueWolf08:

I used the inverse op

BlueWolf08:

would that work?

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 you mean like, since it's (x-12)=12 you added 12? bc yes, that's how you solve these. \(\color{#0cbb34}{\text{End of Quote}}\)

BlueWolf08:

yeah

BlueWolf08:

24x and 24y

mxddi3:

ok so you got the right number. however they are written slightly different. you would write this as x=24 and y=24 make sense?

BlueWolf08:

Oh yes thank you I have a some more if you dont mind helping me

mxddi3:

please post them as another post and someone will be of assistance

BlueWolf08:

Ok